# Galaxy Spin Alignment with Tidal Fields in the SDSS-IV MaNGA Survey

### v1

@@ 3 Spin–Tidal Alignment
The bottom row of Figure 1 shows the spin–tidal alignments for galaxies residing in regions where the tidal anisotropy exceeds the sample median ($q_{\mathrm{med}}=1.235$). The tidal anisotropy, $q$, quantifies the degree of anisotropy in the local tidal field (A. Paranjape et al., 2018) and is defined as
@@
Our analysis, combining a large IFS survey with a directly reconstructed tidal field, provides a more statistically robust detection of the spin–tidal alignment than previous observational studies. The results reveal a mass-dependent trend broadly consistent with recent state-of-the-art hydrodynamic simulations, which have reported an alignment between stellar spins and the cosmic web but only a weak signal for low-mass galaxies (S. Codis et al., 2018; P. Ganeshaiah Veena et al., 2019; K. Kraljic et al., 2020; J. Lee et al., 2021; J.-S. Moon & J. Lee, 2023). We find a stronger alignment for massive galaxies, with a higher detection significance than in lower-mass systems. Nevertheless, we also observe a mass-dependent spin transition, with low-mass galaxies exhibiting a discernible alignment signal, at least in high-$q$ regions.
@@
To evaluate the relative importance of galaxy properties in determining spin alignments, we employ the conditional mutual information (CMI; C. E. Shannon 1948; B. Pandey & S. Sarkar 2017). The CMI, $I(X;Y|Z)$, quantifies how much information $Y$ provides about $X$ beyond what is explained by their mutual correlation with $Z$. When the parameter space spanned by $X$, $Y$, and $Z$ is divided into a total number of $N_{X}\times N_{Y}\times N_{Z}$ bins of small volume, $I(X;Y|Z)$ can be calculated as
@@
where $p(X_{a},Y_{b},Z_{c})$ denotes the probability function of the variables $(X,Y,Z)$ falling within the (a,b,c)(a,b,c)\,th pixel. A higher CMI indicates that $Y$ retains more significant information about $X$ even after controlling for $Z$. It should be noted that the CMI measures statistical correlations across overall distributions rather than systematic trends in the mean value.
Figure 5 plots the CMI, $I(X;Y|Z)$, for $X=|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|$ ($k=1,2,3$), with $Y$ and $Z$ representing various galaxy properties. The values are normalized by the mean and standard deviation derived from 1,000 randomly shuffled samples. The most prominent feature is that $I(X;q|Z)$ and $I(X;\delta|Z)$ exhibit consistently high values, indicating the strongest correlations of spin alignments with environmental factors. The high values of both $I(X;\delta|q)$ and $I(X;q|\delta)$ further imply that these two parameters make independent contributions. In contrast, $I(X;M_{*}|Z)$ remains small, suggesting that stellar mass becomes less influential once environmental effects are considered, as shown in Figure 4. Morphology and environment types also seem to add little information beyond $M_{*}$ or $\delta$. A comparison of the three principal axes shows broadly similar patterns, though alignments with $\hat{\boldsymbol{e}}_{3}$ yield slightly larger CMI, likely reflecting more intricate correlations among galaxy properties.
@@ 4 Summary and Discussion
We have examined the alignment between the spin vectors and the principal axes of the local tidal field using the complete release of the MaNGA IFS survey. The results reveal clear mass-dependent alignment trends. On average, massive galaxies tend to orient their spins perpendicular to the major axis ($\hat{\boldsymbol{e}}_{1}$) and parallel to the intermediate axis ($\hat{\boldsymbol{e}}_{2}$). Low-mass galaxies show the opposite tendency with a transition mass of $M_{*}\sim 10^{10}-10^{10.5}\,M_{\odot}$. This alignment trend becomes particularly pronounced in regions of high tidal anisotropy. No statistically significant alignment signal is detected with respect to the minor axis ($\hat{\boldsymbol{e}}_{3}$). Galaxy morphology affects the alignment strength more than the preferred direction, with S0 and Sa/b galaxies driving the signal in the massive regime. The alignment patterns vary systematically across large-scale environments: cluster galaxies exhibit the strongest alignment, filament galaxies follow similar but weaker trends, and sheet galaxies show a distinctive perpendicular alignment with $\hat{\boldsymbol{e}}_{1}$. Mutual information analysis suggests that tidal anisotropy ($q$) and overdensity ($\rho$) play the primary roles in determining spin alignments.
@@
Various physical mechanisms that influence galaxy spins are expected to imprint distinct alignment patterns with respect to the local tidal field. Specifically, primordial tidal torques are predicted to produce spin alignments along $\hat{\boldsymbol{e}}_{2}$, as described by the linear TTT (e.g., J. Lee & U.-L. Pen, 2000; B. M. Schäfer, 2009). In the nonlinear regime, the vorticity vectors tend to be oriented perpendicular to $\hat{\boldsymbol{e}}_{1}$ (e.g., N. I. Libeskind et al., 2013; J.-S. Moon & J. Lee, 2023), leading galaxy spins to align in the same direction as the local vorticity. Galaxy mergers occur more frequently along $\hat{\boldsymbol{e}}_{3}$ (e.g., E. Tempel & A. Tamm, 2015; E. Tempel et al., 2015), and their orbital angular momenta, being perpendicular to the merger axis, tend to produce spin alignments perpendicular to $\hat{\boldsymbol{e}}_{3}$. In our results, the alignment with $\hat{\boldsymbol{e}}_{2}$ can be naturally linked to the TTT. In particular, massive galaxies in regions of high $q$ and high $\rho$ are likely to experience stronger tidal influences, which can enhance the alignment strength with $\hat{\boldsymbol{e}}_{2}$.
@@
The alignment with $\hat{\boldsymbol{e}}_{1}$, however, is less straightforward to interpret. Since low-mass galaxies are likely to be more strongly affected by nonlinear effects, their observed tendency to align parallel to $\hat{\boldsymbol{e}}_{1}$ seems inconsistent with theoretical expectations. This result also contrasts with results from hydrodynamical simulations such as IllustrisTNG, which predict parallel alignment with $\hat{\boldsymbol{e}}_{1}$ in the massive regime (see, e.g., J. Lee et al., 2021; J.-S. Moon & J. Lee, 2023).
@@
Nevertheless, some observational studies have reported results that cannot be fully explained within this theoretical picture. For instance, Y. Zhang et al. (2015) identified a spin transition with respect to $\hat{\boldsymbol{e}}_{1}$, from parallel to perpendicular, as halo mass increases. They further showed that the alignment with $\hat{\boldsymbol{e}}_{2}$ strengthens with increasing mass and becomes more pronounced in cluster environments, both consistent with our findings. Likewise, J. Lee & J.-S. Moon (2023) found a spin transition among spiral galaxies on void surfaces, from parallel to perpendicular orientations with respect to the direction of maximum compression (i.e., $\hat{\boldsymbol{e}}_{1}$), as stellar mass increases. Such observational deviations may point to the potential influence of baryonic physics, whose role in spin alignments remains poorly understood.
@@
Most IFS-based studies of spin alignment have focused on the orientation of galaxy spins relative to cosmic filaments, revealing various mass-dependent (C. Welker et al., 2020; S. Barsanti et al., 2022) and morphology-dependent (K. Kraljic et al., 2021) spin transitions. Although a direct comparison is difficult, our results show no significant spin alignment signal for $\hat{\boldsymbol{e}}_{3}$, which is somewhat surprising given that filament spines are generally aligned with $\hat{\boldsymbol{e}}_{3}$ (e.g., E. Tempel et al., 2014; N. I. Libeskind et al., 2015). A comparative analysis between spin alignments with filaments and those with tidal fields would be an interesting direction for future work.
@@
This study provides the first statistical investigation of spin–tidal alignment using IFS survey data, revealing clear mass- and environment-dependent alignment trends. The detection of spin alignment with the tidal field highlights the strong influence of large-scale structures in shaping galaxy angular momentum. Our results challenge the conventional picture, indicating that the role of baryonic processes in regulating spin orientations may require further investigation. Future work combining hydrodynamical simulations that incorporate diverse baryonic physics models can help establish a more complete understanding of galaxy spin evolution. Moreover, next-generation IFS surveys with larger samples will enable more detailed statistical analyses of how stellar spin alignments depend on various galaxy properties.
@@
We thank Huiyuan Wang for helpful correspondence regarding the GEMA-VAC data. TO acknowledges support from the Taiwan National Science and Technology Council under Grants Nos. NSTC 112-2112-M-001-034-, NSTC 113-2112-M-001-011- and NSTC 114-2112-M-001-004-, and the Academia Sinica Investigator Project Grant No. AS-IV-114-M03 for the period of 2025–2029. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS website is www.sdss.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics | Harvard & Smithsonian (CfA), the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatório Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.

### v2

@@ 3 Spin–Tidal Alignment
The bottom row of Figure 1 shows the spin–tidal alignments for galaxies residing in regions where the tidal anisotropy exceeds the sample median ($q_{\mathrm{med}}=1.235$). The tidal anisotropy, $q$, quantifies the degree of anisotropy in the local tidal field (2018MNRAS.476.3631P) and is defined as
@@
Our analysis, combining a large IFS survey with a directly reconstructed tidal field, provides a more statistically robust detection of the spin–tidal alignment than previous observational studies. The results reveal a mass-dependent trend broadly consistent with recent state-of-the-art hydrodynamic simulations, which have reported an alignment between stellar spins and the cosmic web but only a weak signal for low-mass galaxies (2018MNRAS.481.4753C; 2019MNRAS.487.1607G; 2020MNRAS.493..362K; 2021ApJ...922....6L; 2023ApJ...945...13M). We find a stronger alignment for massive galaxies, with a higher detection significance than in lower-mass systems. Nevertheless, we also observe a mass-dependent spin transition, with low-mass galaxies exhibiting a discernible alignment signal, at least in high-$q$ regions.
Figure 2: Same as Figure 1, but for lenticulars ($2\leq\mathrm{TType}\leq 0$; top), early-type spirals ($1\leq\mathrm{TType}\leq 3$; middle), and late-type spirals ($4\leq\mathrm{TType}\leq 8$; bottom) in high-$q$ regions.
@@
Figure 3: Same as Figure 1, but for galaxies residing in the sheet (top), filament (middle), and cluster (bottom) environments. Galaxies in void environments are not shown because of their small sample size.
@@
Figure 4: Same as Figure 1, but for galaxies residing in high-$\delta$ (top), intermediate-$\delta$ (middle), and low-$\delta$ (bottom) regions.
Figure 5: Normalized conditional mutual information $I(|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|;Y|Z)$ ($k=1,2,3$) for various combinations of $Y$ and $Z$. Each value is normalized by the mean and standard deviation estimated from 1000 randomly shuffled samples. Darker colors indicate stronger importance of $Y$ in determining spin alignments after controlling for $Z$.
@@
To evaluate the relative importance of galaxy properties in determining spin alignments, we employ the conditional mutual information (CMI; 1948BSTJ...27..379S; 2017MNRAS.467L...6P). The CMI, $I(X;Y|Z)$, quantifies how much information $Y$ provides about $X$ beyond what is explained by their mutual correlation with $Z$. When the parameter space spanned by $X$, $Y$, and $Z$ is divided into a total number of $N_{X}\times N_{Y}\times N_{Z}$ bins of small volume, $I(X;Y|Z)$ can be calculated as
@@
where $p(X_{a},Y_{b},Z_{c})$ denotes the probability function of the variables $(X,Y,Z)$ falling within the $(a,b,c)$th pixel. A higher CMI indicates that $Y$ retains more significant information about $X$ even after controlling for $Z$. It should be noted that the CMI measures statistical correlations across overall distributions rather than systematic trends in the mean value.
Figure 5 plots the CMI, $I(X;Y|Z)$, for $X=|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|$ ($k=1,2,3$), with $Y$ and $Z$ representing various galaxy properties. The values are normalized by the mean and standard deviation derived from 1000 randomly shuffled samples. The most prominent feature is that $I(X;q|Z)$ and $I(X;\delta|Z)$ exhibit consistently high values, indicating the strongest correlations of spin alignments with environmental factors. The high values of both $I(X;\delta|q)$ and $I(X;q|\delta)$ further imply that these two parameters make independent contributions. In contrast, $I(X;M_{*}|Z)$ remains small, suggesting that stellar mass becomes less influential once environmental effects are considered, as shown in Figure 4. Morphology and environment types also seem to add little information beyond $M_{*}$ or $\delta$. A comparison of the three principal axes shows broadly similar patterns, though alignments with $\hat{\boldsymbol{e}}_{3}$ yield slightly larger CMI, likely reflecting more intricate correlations among galaxy properties.
@@ 4 Summary and Discussion
We have examined the alignment between the spin vectors and the principal axes of the local tidal field using the complete release of the MaNGA IFS survey. The results reveal clear mass-dependent alignment trends. On average, massive galaxies tend to orient their spins perpendicular to the major axis ($\hat{\boldsymbol{e}}_{1}$) and parallel to the intermediate axis ($\hat{\boldsymbol{e}}_{2}$). Low-mass galaxies show the opposite tendency with a transition mass of $M_{*}\sim 10^{10}-10^{10.5}\,M_{\odot}$. This alignment trend becomes particularly pronounced in regions of high tidal anisotropy. No statistically significant alignment signal is detected with respect to the minor axis ($\hat{\boldsymbol{e}}_{3}$). Galaxy morphology affects the alignment strength more than the preferred direction, with S0 and Sa/b galaxies driving the signal in the massive regime. The alignment patterns vary systematically across large-scale environments: cluster galaxies exhibit the strongest alignment, filament galaxies follow similar but weaker trends, and sheet galaxies show a distinctive parallel alignment with $\hat{\boldsymbol{e}}_{1}$. Mutual information analysis suggests that tidal anisotropy ($q$) and overdensity ($\rho$) play the primary roles in determining spin alignments.
@@
Various physical mechanisms that influence galaxy spins are expected to imprint distinct alignment patterns with respect to the local tidal field. Specifically, primordial tidal torques are predicted to produce spin alignments along $\hat{\boldsymbol{e}}_{2}$, as described by the linear TTT (e.g., 2000ApJ...532L...5L; 2009IJMPD..18..173S). In the nonlinear regime, the vorticity vectors tend to be oriented perpendicular to $\hat{\boldsymbol{e}}_{1}$ (e.g., 2013ApJ...766L..15L; 2023ApJ...945...13M), leading galaxy spins to align in the same direction as the local vorticity. Galaxy mergers occur more frequently along $\hat{\boldsymbol{e}}_{3}$ (e.g., 2015A&A...576L...5T; 2015MNRAS.450.2727T), and their orbital angular momenta, being perpendicular to the merger axis, tend to produce spin alignments perpendicular to $\hat{\boldsymbol{e}}_{3}$. In our results, the alignment with $\hat{\boldsymbol{e}}_{2}$ can be naturally linked to the TTT. In particular, massive galaxies in regions of high $q$ and high $\rho$ are likely to experience stronger tidal influences, which can enhance the alignment strength with $\hat{\boldsymbol{e}}_{2}$.
@@
The alignment with $\hat{\boldsymbol{e}}_{1}$, however, is less straightforward to interpret. Since low-mass galaxies are likely to be more strongly affected by nonlinear effects, their observed tendency to align parallel to $\hat{\boldsymbol{e}}_{1}$ seems inconsistent with theoretical expectations. This result also contrasts with results from hydrodynamical simulations such as IllustrisTNG, which predict parallel alignment with $\hat{\boldsymbol{e}}_{1}$ in the massive regime (see, e.g., 2021ApJ...922....6L; 2023ApJ...945...13M).
@@
Nevertheless, some observational studies have reported results that cannot be fully explained within this theoretical picture. For instance, 2015ApJ...798...17Z identified a spin transition with respect to $\hat{\boldsymbol{e}}_{1}$, from parallel to perpendicular, as halo mass increases. They further showed that the alignment with $\hat{\boldsymbol{e}}_{2}$ strengthens with increasing mass and becomes more pronounced in cluster environments, both consistent with our findings. Likewise, 2023ApJ...951L..26L found a spin transition among spiral galaxies on void surfaces, from parallel to perpendicular orientations with respect to the direction of maximum compression (i.e., $\hat{\boldsymbol{e}}_{1}$), as stellar mass increases. Such observational deviations may point to the potential influence of baryonic physics, such as anisotropic outflows or heating from stellar and active galactic nucleus feedback (e.g., 2017ApJ...834..169T; 2020MNRAS.492.4268S), whose roles in spin alignments remain poorly understood.
@@
Most IFS-based studies of spin alignment have focused on the orientation of galaxy spins relative to cosmic filaments, revealing various mass-dependent (2020MNRAS.491.2864W; 2022MNRAS.516.3569B) and morphology-dependent (2021MNRAS.504.4626K) spin transitions. Although a direct comparison is difficult, our results show no significant spin alignment signal for $\hat{\boldsymbol{e}}_{3}$, which is somewhat surprising given that filament spines are generally aligned with $\hat{\boldsymbol{e}}_{3}$ (e.g., 2014MNRAS.437L..11T; 2015MNRAS.453L.108L). A comparative analysis between spin alignments with filaments and those with tidal fields would be an interesting direction for future work.
Lastly, we caution that the alignment trends reported in this work are derived from a tidal field smoothed on a scale of 2 $h^{-1}$Mpc. This smoothing scale is comparable to the characteristic size of large-scale structures such as filaments and galaxy clusters, and is therefore a commonly adopted choice (e.g., 2007MNRAS.381...41H; 2007MNRAS.375..489H; 2009ApJ...706..747Z; 2015ApJ...798...17Z; 2017MNRAS.468L.123W). Too small smoothing scales can break the linear approximation of the tidal field, while too large scales can wash out the spin–tidal correlations. However, previous studies have indicated that the detailed trend can vary with the smoothing scale (2012MNRAS.427.3320C; 2013ApJ...762...72T; 2018MNRAS.473.1562W; 2019ApJ...872...37L; 2021ApJ...922....6L; 2023ApJ...945...13M; 2023ApJ...952...82M), suggesting that future analyses using different scales will be valuable for understanding the scale dependence of the observed alignments.
@@
This study provides one of the first observational investigations of spin–tidal alignment using IFS survey data, revealing clear mass- and environment-dependent alignment trends. The detection of spin alignment with the tidal field highlights the strong influence of large-scale structures in shaping galaxy angular momentum. Our results challenge the conventional picture, indicating that the role of baryonic processes in regulating spin orientations may require further investigation. Future work combining hydrodynamical simulations that incorporate diverse baryonic physics models can help establish a more complete understanding of galaxy spin evolution. Moreover, next-generation IFS surveys with larger samples will enable more detailed statistical analyses of how stellar spin alignments depend on various galaxy properties.
@@
We thank the anonymous referee for useful comments that improved the manuscript. We also thank Huiyuan Wang for helpful correspondence regarding the GEMA-VAC data. T.O. acknowledges support from the Taiwan National Science and Technology Council under Grants Nos. NSTC 112-2112-M-001-034-, NSTC 113-2112-M-001-011- and NSTC 114-2112-M-001-004-, and the Academia Sinica Investigator Project Grant No. AS-IV-114-M03 for the period of 2025–2029. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS website is www.sdss.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics | Harvard & Smithsonian (CfA), the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatório Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
@@ Appendix A Gas Spin Alignments
Figure 6: Same as Figure 1, but showing the spin–tidal alignments for the gas spin vectors.
@@
Appendix B Twofold Ambiguity of Spin Vectors
Figure 7: Same as the bottom panels of Figure 1, but recomputed with the sign of $\hat{j}_{r}$ fixed to either ++ (top) or - (bottom).
As discussed in Section 2.1, the twofold ambiguity in the sign of $\hat{j}_{r}$ leads to two possible orientations of a galaxy’s 3D spin vector. Figure 7 tests the impact of this ambiguity by recomputing the spin–tidal alignment in high-$q$ regions (same as the bottom panels of Figure 1) with all signs of $\hat{j}_{r}$ fixed to either ++ (radially outward) or - (radially inward). Both cases reproduce the same qualitative alignment trends as the fiducial analysis, showing only a modest reduction in the detection significance.
Galaxies whose inferred 3D spin orientation changes substantially under different sign choices tend to be less reliable. This consideration motivates our use of the averaged orientation, which provides a more stable ensemble statistic. For this reason, we adopt the averaged orientation as our fiducial choice, while noting that our main conclusions are qualitatively insensitive to the specific sign convention.